1 Feb. 2024 LOW NOISE 3 (~45-40 MpC) >>>> UPS switch off >>>> 11:05 UTC - switched off pumps and sensors at 600N 11:10 UTC - switched off pumps and sensors at 600W (note: turbo pumps take a while, also 1hour) to come to rest) checked power consumption (regime) is about 500W 13:38 UTC 600N rack off 13:53 UTC 600W rack off 14:57 UTC 600N UPS off 15:06 UTC 600W UPS off 15:20 UTC (approx) 600N and 600W back to standard condition Observe largest effect on ENV_CEB_MAG_V, 12.8 Hz sidebands of 50Hz, 100Hz and 150Hz. >>>> HVAC switch off >>>>>>>> NEB 16:01 AHU off 16:02 glitch hrec WEB 16:01 AHU off 16:06 chillers off 16:11 cold water pump off 17:23 (~) back on CEB 16:14 clean-rooms AHU OFF 16:19 hall ahu off 16:24 det AHU off 16:29 inj AHU off 17:00 inj AHU on 17:05 det AHU on 17:10 ceb AHU on 17:14 cr AHU on 16:29 glitch hrec 16:56:16 glitch hrec 16:59:50 ITF unlock. 17:13 NEB on 17:23 WEB on disconnected NI and WI ring heaters power supplies 17:45:30 relocked, but very large noise below 100 Hz 17:49:30 hrec glitch 18:16 hrec glitch Fede connects to mass all cables (observed no effect on glitches: still there :-(((( Then reconnect back all RH power supplies 19:22:30 TF relocked about 19:00 UTC CEB hall HVAC: set in manual mode and opened cold water valves >>> hall temperature decreses.... about -2deg in 45 minutes 19:35:30 hrec glitch (expected at ~19:38 with estimated rate of 27minutes) 19:47 ITF unlocks 20:24:30 ITF locked in low noise 3